OpenStudy (anonymous):

the parabolas y=-3x^2+10x-6 and y=-3x^2-17x+2 intersect at A.they do not intersect B.x=sq rt2/3 and x=-sq rt2/3 C.x=8/27 D.x=8/27 and x=-4/27

OpenStudy (anonymous):

your job is to set \[-3x^2+10x-6=-3x^2-17x+2\]and solve for \(x\)

OpenStudy (anonymous):

shouldn't be too hard since it is the same as \[10x-6=-17x+2\]

OpenStudy (anonymous):

a) add \(17x\) to both sides b) add \(6\) to both sides c) divide both sides by \(27\)

OpenStudy (anonymous):

whats the answer i got c

OpenStudy (anonymous):

@mathmate can u help ,e

OpenStudy (mathmate):

What did you get for x?

OpenStudy (anonymous):

17

OpenStudy (mathmate):

How did you get 17? Did you solve for x as @satellite73 showed you?

OpenStudy (anonymous):

yea i got C is that correct

OpenStudy (mathmate):

What did you get for x again?

OpenStudy (anonymous):

17

OpenStudy (mathmate):

C doesn't say x=17! :)

OpenStudy (anonymous):

none of them do so would it be A

OpenStudy (mathmate):

@Twilliams250 There are two ways to do math: 1. Get the answers, and get 90 for assignments, but fail at exams. 2. Put work in there, learn how to solve problems, get perhaps 70 for assignments, but get the same for exams. Your choice is:?

OpenStudy (anonymous):

2

OpenStudy (mathmate):

Very good! So you want to follow what @satellite73 gave you as instructions to solve for the value of x. Then come back here and check you answers. If the answer is not correct, we'll help you get there.

OpenStudy (anonymous):

need help

OpenStudy (mathmate):

Start with @satellite73's post: shouldn't be too hard since it is the same as 10x−6=−17x+2 a) add 17x to both sides b) add 6 to both sides c) divide both sides by 27 Can you do them step by step?

OpenStudy (anonymous):

still having trouble

OpenStudy (mathmate):

Can you reproduce the starting statement, meaning the equation "10x-6=-17x+2"

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Now below it, work out step (a) and tell me what you've got.